OpenStudy (anonymous):

A garden table and a bench cost $752 combined. The garden table costs $98 less than the bench. What is the cost of the bench?

OpenStudy (anonymous):

garden table cost 278 and bench cost 474

OpenStudy (anonymous):

both are wrong

OpenStudy (anonymous):

$752/2 = 376 376-96 = 280 376+96 =472 therefore the bench costs 472?

OpenStudy (anonymous):

98 not 96 problem say 98

OpenStudy (anonymous):

so its 474

OpenStudy (anonymous):

ah yeah that is what i would have thought

OpenStudy (anonymous):

i did it same way u did it rahulp but its 98 not 96